OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (shane_b):

You need to do a system of equations to solve this: \[0.25x+0.10y=$7.00\]\[x+y=46\] Can you solve it from here?

OpenStudy (anonymous):

this is were i got stuck

OpenStudy (shane_b):

First, you need to you rearrange one of the equations so it looks like x=... or y=... The second equation is simpler so can you put it in terms of x or y?

OpenStudy (anonymous):

ok

OpenStudy (shane_b):

To nudge you along: \[x+y=46\]...what can you do to make it look like x=??? ?

OpenStudy (anonymous):

im confused sorry

OpenStudy (shane_b):

Can you rearrange x+y=46 so that x is by itself on one side of the equation?

OpenStudy (shane_b):

You need to subtract y from both sides...which will leave you with?

OpenStudy (anonymous):

10

OpenStudy (shane_b):

x+y=46...and you need to subtract y from both sides to isolate x by itself. How in the world do you get 10?

OpenStudy (shane_b):

@Luis_Rivera: What problem is that for? It doesn't look like it's this one!

OpenStudy (shane_b):

@niyabae: Obviously you're not getting this and I don't have a lot of time to teach you basic algebraic manipulations. The answer is 16 quarters, 30 dimes. Good luck!

OpenStudy (shane_b):

Just so you know how I got those values:\[.25x+.10y=$7.00\]\[x+y=46\]\[x=46-y\]\[.25(46-y)+.10y=$.700\]\[**x=16\]\[.25(16)+.10y=$7.00\]\[**y=30\]